# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ Abstract.
This paper studies Positivstellensätze for a polynomial matrix subject to polynomial matrix inequality constraints with universal quantifiers. We first present a Scherer-Hol-type Positivstellensatz under the Archimedean condition. When the objective is a scalar polynomial, we further provide a sparse Scherer-Hol-type Positivstellensatz in the presence of correlative sparsity. Next, without assuming the Archimedean condition, we derive Putinar-Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze under the same setting. These results can be viewed as common generalizations of corresponding Positivstellensätze in the cases of polynomials, polynomials with universal quantifiers, and polynomial matrices. For the proofs, techniques from $\star$-algebra, real algebraic geometry, operator theory, and convex optimization are employed. Applications of the established Positivstellensätze to robust polynomial matrix optimization are also discussed.
@@ 1. Introduction
Positivstellensätze are fundamental results in real algebraic geometry, asserting under which conditions a polynomial is guaranteed to be positive on a given set [6, 31, 35, 44]. Most Positivstellensätze achieve this by expressing the polynomial using sums of squares (SOS). These powerful results offer constructive methods to certify whether a polynomial $f$ is positive over a basic semialgebraic set defined by
@@
and have widespread applications in areas like optimization, algebraic geometry, control theory, and more where polynomial positivity is a key concern [5, 14, 29, 31, 36].
Depending on whether the semialgebraic set $\mathbf{K}$ is assumed to be compact, Positivstellensätze are divided into two categories. When $\mathbf{K}$ is a compact polyhedron with non-empty interior, Handelman’s Positivstellensatz [13] states that a polynomial $f$ that is positive on $\mathbf{K}$ can be expressed as a positive linear combination of cross-products of $g_{j}$’s. In the case of a more general compact basic semialgebraic set $\mathbf{K}$, Schmüdgen [47] demonstrated that if a polynomial $f$ is positive on $\mathbf{K}$, then it belongs to the preordering generated by $g_{j}$’s, that is, $f$ can be represented as an SOS-weighted combination of cross-products of $g_{j}$’s. Putinar’s Positivstellensatz [41] provides an alternative representation that avoids the need for cross-products of $g_{j}$’s under the Archimedean condition (slightly stronger than compactness). Building on Putinar’s Positivstellensatz, Lasserre [26, 28, 29] introduced a moment-SOS hierarchy of semidefinite relaxations for polynomial optimization. This framework generates a non-decreasing sequence of lower bounds converging to the optimal value of a given polynomial optimization problem. When $f$ is only nonnegative on $\mathbf{K}$, Lasserre and Netzer [30] provided an SOS approximation of $f$ via high-degree perturbations.
There have been several attempts to provide certificates of positivity of a polynomial over a non-compact semialgebraic set. Without requiring compactness of $\mathbf{K}$, Krivine-Stengle Positivstellensatz [25, 50] states that a polynomial $f$ is positive over $\mathbf{K}$ if and only if $\psi f=1+\phi$ for some $\psi,\phi$ from the preordering generated by $g_{j}$’s. Pólya [38] demonstrated that if $f$ is homogeneous and positive on +${\mathbb{R}}^{n}_{+}\setminus\{\mathbf{0}\}$, then multiplying $f$ by some power of $\sum_{i=1}^{n}x_{i}$, one obtains a polynomial with nonnegative coefficients. Dickinson and Povh [9] generalized Pólya’s Positivstellensatz for homogeneous polynomials being positive on the intersection +${\mathbb{R}}^{n}_{+}\cap\mathbf{K}\setminus\{\mathbf{0}\}$, where $g_{j}$’s are assumed to be also homogeneous. Reznick [43] proved that after be multiplying by some power of $\sum_{i=1}^{n}x_{i}^{2}$, any positive definite (PD) form is a sum of even powers of linear forms. Putinar and Vasilescu [42] extended Reznick’s result to the constrained case where $f$ and $g_{j}$’s are homogeneous polynomial of even degree. When $\mathbf{K}$ is non-compact, SOS-structured certificates of positivity of $f$ over $\mathbf{K}$ could be also established by investigating various geometric objects associated with the data, e.g., gradient varieties [37], principal gradient tentacles [49], truncated tangency varieties [52], and the polar varieties [11].
@@
where each $\sigma_{j}$, $j=0,1,\ldots,s$, are SOS polynomials. They also investigated the corresponding moment problem on the semialgebraic set $\mathcal{U}$. As an important application, their results could be used to solve semi-infinite optimization problems which are highly challenge.
Most Positivstellensätze could be generalized to the matrix setting that both $f$ and $g_{j}$’s are polynomial matrices. Scherer and Hol [45] developed a matrix-version of Putinar’s Positivstellensatz. Cimprič [7] extended Krivine-Stengle Positivstellensatz to the case of polynomial matrices with polynomial constraints. A matrix version of Handelman’s Positivstellensatz was proposed in [32]. Building on Scherer and Hol’s Positivstellensatz, Dinh et al. [10] generalized the classical Schmüdgen, Putinar-Vasilescu, and Dickinson–Povh Positivstellensätze to the polynomial matrix setting. Given the projections of two semialgebraic sets defined by polynomial matrix inequalities (PMI), Klep and Nie [22] provided a matrix Positivstellensatz with lifting polynomials to determine whether one is contained in the other. These generalizations have a wide range of applications, particularly in areas such as optimal control, systems theory [15, 16, 19, 40, 51].
Computing SOS-structured representations involved in Positivstellensätze could be typically cast as semidefinite programs (SDP). However, the size of the SDPs grows rapidly with the problem dimension. Hence from the perspective of computation, it becomes appealing to develop sparse versions of Positivstellensätze for sparse data, e.g., correlative sparsity [21, 27, 53], term sparsity [21, 33, 54, 55, 56], and matrix (chordal) sparsity [57, 58].
Due to estimation errors or lack of information, the data of real-world problems often involve uncertainty. As a result, ensuring the robustness of PMIs over a prescribed set with uncertainty is crucial for some safety-critical applications with little tolerance for failure [3]. Consequently, Positivstellensätze for polynomial matrices with UQs will be a powerful mathematical tool for addressing this issue, which serves as the primary motivation for this work. Precisely, consider the semialgebraic set
@@
Lasserre-Netzer Lasserre and Netzer [30] Theorem 4.4 Theorem 4.4 Theorem 4.4
@@
Contributions. We generalize several classical Positivstellensätze for scalar polynomials to the matrix setting with UQs (see Tables 1 and 2 for summaries). In the following, we highlight the main results of this paper. Throughout the paper, let $\nu$ be a fixed Borel measure on ${\mathbb{R}}^{m}$ with support $\mathrm{supp}(\nu)=\mathcal{Y}$ and satisfying $\int_{\mathcal{Y}}|h({\boldsymbol{y}})|\mathrm{d}\nu({\boldsymbol{y}})<\infty$ for all $h({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]$. For any $H({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{p}$, let us write $H({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\beta}}\in\mbox{supp}_{{\boldsymbol{y}}}(H)}H_{{\boldsymbol{\beta}}}({\boldsymbol{x}}){\boldsymbol{y}}^{{\boldsymbol{\beta}}}$ as a polynomial matrix in ${\boldsymbol{y}}$ with coefficient matrices $H_{{\boldsymbol{\beta}}}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and let
@@
The main foundational result of this paper is the following Scherer-Hol-type Positivestellensatz which holds under the Archimedean condition (Assumption 3.1) and the Carleman condition on $\nu$ (Assumption 3.2).
@@
Theorem A. (Theorem 3.1) Suppose that Assumptions 3.1 and 3.2 hold. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}$, then there exists SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ such that
$F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
See Section 2.1 for the definition of the product $\langle\cdot,\cdot\rangle_{p}$. We adopt a notably nontrivial operator-theoretic approach to prove Theorem A, utilizing techniques such as the Gelfand-Naimark-Segal (GNS) construction for a carefully constructed $\star$-algebra, and the spectral theory of self-adjoint operators in a Hilbert space.
@@
Then we consider the particular case of $p=1$, namely, the objective is a scalar polynomial. Suppose that $\mathcal{X}$ is defined by multiple PMIs with UQs ($G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,j=1,\ldots,s$). We assume the presence of correlative sparsity in the problem data, which implies that the variables ${\boldsymbol{x}}$ decompose as a union of subsets ${\boldsymbol{x}}=\cup_{\ell=1}^{t}{\boldsymbol{x}}(\mathcal{I}_{\ell})$ such that $\{G_{j}\}_{j=1}^{s}=\cup_{\ell=1}^{t}\{G_{j}\}_{j\in\mathcal{J}_{\ell}}$ and $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{q_{j}}$, $j\in\mathcal{J}_{\ell}$, $\ell=1,\ldots,t$ (see Assumption 3.3). Under these conditions, we could give the following sparse version of the Scherer-Hol-type Positivstellensatz.
@@
Theorem B. (Theorem 3.2) Suppose that $\mathcal{Y}$ is compact, Assumption 3.3 holds for $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $\mathcal{X}$, and Assumption 3.1 holds with respect to each ${\boldsymbol{x}}(\mathcal{I}_{\ell})$. If $f>0$ on $\mathcal{X}$, then there exist SOS polynomials $\sigma_{\ell,0}\in{\mathbb{R}}[{\boldsymbol{x}}],\sigma_{\ell,j}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]$ such that
@@
$f({\boldsymbol{x}})=\sum_{\ell=1}^{t}\left(\sigma_{\ell,0}({\boldsymbol{x}})+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\langle\sigma_{\ell,j}({\boldsymbol{x}},{\boldsymbol{y}}),G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\rangle\mathrm{d}\nu({\boldsymbol{y}})\right).$
@@
Next, building on Theorem A and existing techniques, we establish several Positivstellensätze for polynomial matrices with UQs and without assuming the Archimedean condition. The first such Positivstellensatz is of Putinar–Vasilescu-type.
@@
Theorem C. (Theorem 4.1) Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$ of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and $\deg\Sigma_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}\Sigma=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The second such Positivstellensatz is of Pólya-type.
@@
Theorem D. (Theorem 4.2) Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{{\boldsymbol{\alpha}}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{{\boldsymbol{\alpha}}}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy each $S_{0,{\boldsymbol{\alpha}}}\succ 0$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all ${\boldsymbol{y}}\in\mathcal{Y}$, such that
$\left(\sum_{i=1}^{n}x_{i}\right)^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The third such Positivstellensatz is of Lasserre-Netzer-type.
Theorem E. (Theorem 4.4) Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is positive semidefinite (PSD) on $\mathcal{X}\cap[-1,1]^{n}$. Then for any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1},k\geq\ell_{2}$, it holds
$F({\boldsymbol{x}})+\varepsilon\left(1+\sum_{i=1}^{n}x_{i}^{2d}\right)I_{p}=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg(\Sigma_{0})\leq 2d$, $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$, and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$.
To prove Theorem E, we studied the exponentially bounded matrix-valued moment problem, providing a characterization for an exponentially bounded matrix-valued sequence to admit a matrix-valued representing measure supported on the set $\mathcal{X}\cap[-C,C]^{n}$ for a given $C>0$ (Theorem 4.3). This result extends the work of Berg and Maserick [2] to the matrix case and offers novel insights on its own.
@@
The rest of this paper is organized as follows. In Section 2, we review some preliminary concepts. Section 3 presents a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs under the Archimedean condition, along with a sparse version in the presence of correlative sparsity. In Section 4, without relying on the Archimedean condition, we derive Putinar–Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze for polynomial matrices with UQs. Section 5 briefly discusses applications of the established Positivstellensätze to robust polynomial matrix optimization. For the sake of readability, some lengthy and technical proofs are deferred to Section 6. Conclusions are given in Section 7.

### v4

@@ Abstract.
This paper investigates Positivstellensätze for polynomial matrices subject to universally quantified polynomial matrix inequality constraints. We first establish a matrix-valued Positivstellensatz under the Archimedean condition, incorporating universal quantifiers. For scalar-valued polynomial objectives, we further develop a sparse Positivstellensatz that leverages correlative sparsity patterns within these quantified constraints. Moving beyond the Archimedean framework, we then derive two generalized Positivstellensätze under analogous settings. These results collectively unify and extend foundational theorems in three distinct contexts: classical polynomial Positivstellensätze, their universally quantified counterparts, and matrix polynomial formulations. Applications of the established Positivstellensätze to robust polynomial matrix inequality constrained optimization are also investigated.
@@ 1. Introduction
Positivstellensätze are fundamental results in real algebraic geometry, asserting under which conditions a polynomial is guaranteed to be positive on a given set [5, 26, 31, 41]. Most Positivstellensätze achieve this by expressing the polynomial using sums of squares (SOS). These powerful results offer constructive methods for certifying whether a polynomial $f$ is positive over a basic semialgebraic set defined by polynomial inequalities
@@
and have widespread applications in areas such as optimization, algebraic geometry, control theory, and more where polynomial positivity is a key concern [4, 14, 25, 26, 34].
Depending on whether the semialgebraic set $\mathbf{K}$ is assumed to be compact, Positivstellensätze are divided into two categories. When $\mathbf{K}$ is a compact polyhedron with non-empty interior, Handelman’s Positivstellensatz [12] states that a polynomial $f$ that is positive on $\mathbf{K}$ can be expressed as a positive linear combination of cross-products of $g_{j}$’s. In the case of a more general compact basic semialgebraic set $\mathbf{K}$, Schmüdgen [43] demonstrated that if a polynomial $f$ is positive on $\mathbf{K}$, then it belongs to the preorder generated by $g_{j}$’s, that is, $f$ can be represented as an SOS-weighted combination of cross-products of $g_{j}$’s. Putinar’s Positivstellensatz [39] provides an alternative representation that avoids the need for cross-products of $g_{j}$’s under the Archimedean condition (slightly stronger than compactness). Building on Putinar’s Positivstellensatz, Lasserre [22, 24, 25] introduced a moment-SOS hierarchy of semidefinite relaxations for polynomial optimization. This framework generates a non-decreasing sequence of lower bounds converging to the optimal value of a given polynomial optimization problem.
There have been several attempts to provide certificates of positivity of a polynomial over a non-compact semialgebraic set. Without requiring compactness of $\mathbf{K}$, Krivine-Stengle Positivstellensatz [21, 46] states that a polynomial $f$ is positive over $\mathbf{K}$ if and only if $\psi f=1+\phi$ for some $\psi,\phi$ from the preorder generated by $g_{j}$’s. Pólya [35] demonstrated that if $f$ is homogeneous and positive on ${\mathbb{R}}^{n}_{+}\setminus\{\mathbf{0}\}$, then multiplying $f$ by some power of $\sum_{i=1}^{n}x_{i}$, one obtains a polynomial with nonnegative coefficients. Dickinson and Povh [7] generalized Pólya’s Positivstellensatz for homogeneous polynomials being positive on the intersection ${\mathbb{R}}^{n}_{+}\cap\mathbf{K}\setminus\{\mathbf{0}\}$, where $g_{j}$’s are assumed to be also homogeneous. Reznick [40] proved that, after multiplying by some power of $\sum_{i=1}^{n}x_{i}^{2}$, any positive definite (PD) form is a sum of even powers of linear forms. Putinar and Vasilescu [38] extended Reznick’s result to the constrained case where $f$ and $g_{j}$’s are homogeneous polynomials of even degree. When $\mathbf{K}$ is non-compact, SOS-structured positivity certificates of $f$ over $\mathbf{K}$ could be also established by investigating various geometric objects associated with the data, e.g., gradient varieties [33], principal gradient tentacles [45], truncated tangency varieties [48], and the polar varieties [10].
@@
where each $\sigma_{j}$, $j=0,1,\ldots,s$, are SOS polynomials. They also investigated the corresponding moment problem on the semialgebraic set $\mathcal{U}$. As an important application, their results could be used to solve semi-infinite optimization problems which are highly challenging.
Most Positivstellensätze could be generalized to the matrix setting that both $f$ and $g_{j}$’s are polynomial matrices. Scherer and Hol [42] developed a matrix-version of Putinar’s Positivstellensatz. Cimprič [6] extended the Krivine-Stengle Positivstellensatz to the case of polynomial matrices with polynomial constraints. A matrix version of Handelman’s Positivstellensatz was proposed in [27]. Building on Scherer and Hol’s Positivstellensatz, Dinh et al. [8] generalized the classical Schmüdgen, Putinar-Vasilescu, and Dickinson–Povh Positivstellensätze to the polynomial matrix setting. Given the projections of two semialgebraic sets defined by polynomial matrix inequalities (PMI), Klep and Nie [19] provided a matrix Positivstellensatz with lifting polynomials to determine whether one is contained in the other. These generalizations have a wide range of applications, particularly in areas such as optimal control, systems theory [13, 15, 17, 37, 47].
Computing SOS-structured representations involved in Positivstellensätze could be typically cast as semidefinite programs (SDP). However, the size of the SDPs grows rapidly with the dimension of the problem. Hence, from the perspective of computation, it becomes appealing to develop sparse versions of Positivstellensätze for sparse data, e.g., correlative sparsity [32, 23, 49], term sparsity [32, 29, 51, 52, 50], and matrix (chordal) sparsity [54, 53].
Due to estimation errors or lack of information, the data from real-world problems often involve uncertainty. As a result, ensuring the robustness of PMIs over a prescribed set with uncertainty is crucial for some safety-critical applications with little tolerance for failure [3]. Consequently, Positivstellensätze for polynomial matrices with UQs will be a powerful mathematical tool for addressing this issue, which serves as the primary motivation for this work. Precisely, consider the semialgebraic set
@@
Contributions. We generalize several classical Positivstellensätze from scalar polynomials to the matrix setting with UQs (see Tables 1 and 2 for summaries). In the following, we highlight the main results of this paper. Throughout the paper, let $\nu$ be a fixed Borel measure on ${\mathbb{R}}^{m}$ with support $\mathrm{supp}(\nu)=\mathcal{Y}$ and satisfying $\int_{\mathcal{Y}}|h({\boldsymbol{y}})|\mathrm{d}\nu({\boldsymbol{y}})<\infty$ for all $h({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]$. For any $H({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{p}$, let us write $H({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\beta}}\in\mbox{supp}_{{\boldsymbol{y}}}(H)}H_{{\boldsymbol{\beta}}}({\boldsymbol{x}}){\boldsymbol{y}}^{{\boldsymbol{\beta}}}$ as a polynomial matrix in $\boldsymbol{y}$} with coefficient matrices $H_{{\boldsymbol{\beta}}}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and let
@@
The main result of this paper is the following matrix-valued Positivestellensatz incorporating universal quantifiers, which holds under the Archimedean condition (Assumption 3.1) and the Carleman condition on $\nu$ (Assumption 3.2). See Section 2 for the definition of the product $\langle\cdot,\cdot\rangle_{p}$.
@@
Theorem A. (Theorem 3.1) Suppose that Assumptions 3.1 and 3.2 hold. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}$, then there exist SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ such that
$F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
Then we consider the particular case of $p=1$, namely, the objective is a scalar polynomial. Suppose that $\mathcal{X}$ is defined by multiple PMIs with UQs ($G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,j=1,\ldots,s$). We assume the presence of correlative sparsity in the problem data, which implies that the variables $\boldsymbol{x}$} decompose as a union of subsets ${\boldsymbol{x}}=\cup_{\ell=1}^{t}{\boldsymbol{x}}(\mathcal{I}_{\ell})$ such that $\{G_{j}\}_{j=1}^{s}=\cup_{\ell=1}^{t}\{G_{j}\}_{j\in\mathcal{J}_{\ell}}$ and $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{q_{j}}$, $j\in\mathcal{J}_{\ell}$, $\ell=1,\ldots,t$ (see Assumption 3.3). Under these conditions, we could give the following sparse Positivstellensatz that leverages correlative sparsity patterns within these quantified constraints.
@@
Theorem B. (Theorem 3.2) Suppose that $\mathcal{Y}$ is compact, Assumption 3.3 holds for $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $\mathcal{X}$, and Assumption 3.1 holds with respect to each ${\boldsymbol{x}}(\mathcal{I}_{\ell})$. If $f>0$ on $\mathcal{X}$, then there exist SOS polynomials $\sigma_{\ell,0}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$, and SOS polynomial matrices $S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{q_{j}}$, $j\in\mathcal{J}_{\ell}$, such that
@@
$f({\boldsymbol{x}})=\sum_{\ell=1}^{t}\left(\sigma_{\ell,0}({\boldsymbol{x}}(\mathcal{I}_{\ell}))+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\langle S_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}),G_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}})\rangle\mathrm{d}\nu({\boldsymbol{y}})\right).$
@@
Next, building on Theorem A and existing techniques, we establish two generalized Positivstellensätze for polynomial matrices with UQs and without assuming the Archimedean condition. The first is a Positivstellensatz for non-compact case.
@@
Theorem C. (Theorem 4.1) Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$} of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and $\deg S_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}S=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The second is a Positivstellensatz on the nonnegative orthant.
@@
Theorem D. (Theorem 4.2) Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $P_{0}=\sum_{{\boldsymbol{\alpha}}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $P=\sum_{{\boldsymbol{\alpha}}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and satisfy each $P_{0,{\boldsymbol{\alpha}}}\succ 0$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all ${\boldsymbol{y}}\in\mathcal{Y}$, such that
$\left(\sum_{i=1}^{n}x_{i}\right)^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The inhomogeneous counterparts of Theorems C and D appear in Corollaries 4.1 and 4.3, respectively.
As applications, leveraging the SOS-structured certificates provided by the established Positivstellensätze, we propose hierarchies of SDP relaxations for the robust PMI constrained optimization problem and analyze their convergence.
@@
The rest of this paper is organized as follows. In Section 2, we review some preliminary concepts. Section 3 presents a matrix-valued Positivstellensatz for polynomial matrices with UQs under the Archimedean condition, along with a sparse version in the presence of correlative sparsity. In Section 4, without relying on the Archimedean condition, we derive two generalized Positivstellensätze for polynomial matrices with UQs. In Section 5, we apply the established Positivstellensätze to robust PMI constrained optimization. For the sake of readability, some lengthy and technical proofs are deferred to Section 6. Conclusions are given in Section 7.